This is a two part question.Suppose that all firms in a perfectly competitive market are identical and have the following cost function C(Q)= 16Q with MC
Answer #1

(Ca)= 16 ton Mc = 20 Ace 16 so JAC Al is minimized when och AC- 164 - 8 ФрдчP-зо че When P=8 00=80-32= 48 no. of firin - do 4